# Robust Control Lyapunov-Value Functions for Nonlinear Disturbed Systems

### v2

@@ III-C Existence of R-CLVF
The R-CLVF exists on an open set $\mathcal{D}_{\gamma}$ (or $\mathbb{R}^{n}$) if the system is exponentially stabilizable (under rate $\gamma$) to its SRCIS from $\mathcal{D}_{\text{ROES}}$ (or $\mathbb{R}^{n}$), despite worst-case disturbance. Further $\mathcal{D}_{\gamma}=\mathcal{D}_{\text{ROES}}$.
@@
Assume the system is exponentially stabilizable to the SRCIS. Using the Definition 2, we have $\forall\lambda\in\Lambda$, $\exists u^{*}(\cdot)\in\mathbb{U}$, $\exists c>0$ s.t.
Plug in equation (2),
Plug in $\ell(x)=h(x)-V^{\infty}_{m}=\|x\|-V^{\infty}_{m}$, we have
$\ell(\xi(s;t,x,u^{*}(\cdot),\lambda[u^{*}]))$
$=$ $h(\xi(s;t,x,u^{*}(\cdot),\lambda[u^{*}]))-V^{\infty}_{m}$
$=$ $\|\xi(s;t,x,u^{*}(\cdot),\lambda[u^{*}])\|-\max_{a\in\partial\mathcal{I}_{m}}\|a\|$
and
$\|\xi(s;t,x,u^{*}(\cdot),\lambda[u^{*}])\|-\min_{a\in\partial\mathcal{I}_{m}}\|a\|$
$=$ $\min_{a\in\partial\mathcal{I}_{m}}\bigl{(}\|\xi(s;t,x,u^{*}(\cdot),\lambda[u^{*}])\|-\|a\|\bigl{)}$
$\leq$ $\min_{a\in\partial\mathcal{I}_{m}}\bigl{(}\|\xi(s;t,x,u^{*}(\cdot),\lambda[u^{*}])-a\|\bigl{)}$
$\leq$ $ce^{-\gamma(s-t)}\min_{a\in\partial\mathcal{I}_{m}}\|x-a\|,$ (33)
where we used equation (5) for the last inequality. Multiply $e^{\gamma(s-t)}$ on both side
which holds for all $s\in[t,0]$. Therefore
$V_{\gamma}(x,t)=$ $\max_{s\in[t,0]}e^{\gamma(s-t)}\left(\ell(\xi(s;t,x,u^{*}(\cdot),\lambda[u^{*}]))\right)$
$\leq$ $c\min_{a\in\partial\mathcal{I}_{m}}\|x-a\|.$
This upper bound $c\min_{a\in\partial\mathcal{I}_{m}}\|x-a\|$ is independent of $t$, therefore as $t\rightarrow-\infty$, we have $V_{\gamma}^{\infty}(x)\leq c\min_{a\in\partial\mathcal{I}_{m}}\|x-a\|$. Since the R-CLVF monotonically increases, we conclude that the limit in (14) exists $\forall x\in\mathcal{D}_{\text{ROES}}$, and $\mathcal{D}_{\gamma}=\mathcal{D}_{\text{ROES}}$.
This Lemma shows that for nonlinear systems, the existence of the R-CLVF can be justified by analyzing the system dynamics.
One natural question to ask is, does the existence of the R-CLVF in turn imply the exponential stabilizability of the system? The answer is yes, as provided in the following Lemma.
Lemma 6
The system can be exponentially stabilized to its smallest robustly control invariant set $\mathcal{I}_{m}$ from $\mathcal{D}_{\gamma}\setminus\mathcal{I}_{m}$ (or $\mathbb{R}^{n}\setminus\mathcal{I}_{m}$), if the R-CLVF exists in $\mathcal{D}_{\gamma}$ (or $\mathbb{R}^{n}$).
Proof:
Assume the limit in (14) exists in $\mathcal{D}_{\gamma}$. For any initial state $x\in\mathcal{D}_{\gamma}\setminus\mathcal{I}_{m}$, consider the optimal trajectory $\xi(s;t,x,u^{*}(\cdot),\lambda[u^{*}])$ $\forall t\leq s\leq 0$. From Proposition 5:
$D_{x}V_{\gamma}^{\infty}(x)\cdot f(x,u^{*},d^{*})=\dot{V}_{\gamma}^{\infty}\leq-\gamma V^{\infty}_{\gamma}.$
@@
$V_{\gamma}^{\infty}\big{(}\xi(s;t,x,u^{*}(\cdot),\lambda[u^{*}])\big{)}\leq e^{-\gamma(s-t)}V_{\gamma}^{\infty}(x).$ (34)
Since $V_{\gamma}(x,0)\leq V_{\gamma}^{\infty}(x)$, we have:
$\|\xi(s;t,x,u^{*}(\cdot),\lambda[u^{*}])-p\|$
$\leq$ $V_{\gamma}^{\infty}\big{(}\xi(s;t,x,u^{*}(\cdot),\lambda[u^{*}])\big{)}+V^{\infty}_{m}.$
Therefore, $\forall s\in[t,0]$ we have:
$=$ $\min_{a\in\partial\mathcal{I}_{m}}\|\xi(s;t,x,u^{*}(\cdot),\lambda[u^{*}])-p+p-a\|$
$\leq$ $\|\xi(s;t,x,u^{*}(\cdot),\lambda[u^{*}])-p\|+\min_{a\in\partial\mathcal{I}_{m}}\|p-a\|$
$\leq$ $\|\xi(s;t,x,u^{*}(\cdot),\lambda[u^{*}])-p\|+V^{\infty}_{m}$
$\leq$ $V_{\gamma}^{\infty}\big{(}\xi(s;t,x,u^{*}(\cdot),\lambda[u^{*}])\big{)}+2V^{\infty}_{m}.$
Plugging in (34) gives us
$\leq$ $e^{-\gamma(s-t)}V_{\gamma}^{\infty}(x)+2V^{\infty}_{m}$
$\leq$ $e^{-\gamma(s-t)}k_{1}\min_{a\in\partial\mathcal{I}_{m}}\|x-a\|+e^{-\gamma(s-t)}k_{2}\min_{a\in\partial\mathcal{I}_{m}}\|x-a\|$
$=$ $e^{-\gamma(s-t)}(k_{1}+k_{2})\min_{a\in\partial\mathcal{I}_{m}}\|x-a\|.$
where
$k_{1}=\frac{V_{\gamma}^{\infty}(x)}{\min_{a\in\partial\mathcal{I}_{m}}\|x-a\|},\hskip 10.00002ptk_{2}=\frac{2V^{\infty}_{m}}{e^{\gamma t}\min_{a\in\partial\mathcal{I}_{m}}\|x-a\|},$
and $0<k_{1},k_{2}<\infty$ for any given $x\notin\mathcal{I}_{m}$. In other words, the controlled system can be locally exponentially stabilized to $\mathcal{I}_{m}$ from $\mathcal{D}_{\text{ROES}}$, if the R-CLVF exists on $\mathcal{D}_{\gamma}$. Further, if the R-CLVF exists on $\mathbb{R}^{n}$, the above result holds globally.
@@
III-D Impact of the Gamma
We highlight the impact of $\gamma$ in this part. The first result is that for all $\gamma\geq 0$, the zero sub-level sets of the R-CLVFs are all the same, presented in Lemma 8. This Lemma guarantees that the shape and size of SRCIS only depend on dynamics and the norm we pick, and for all different $\gamma$, we stabilize the system to the same SRCIS. Denote the zero sub-level sets of TV-R-CLVF and R-CLVF as
[…]
The TV-R-CLVF (and the R-CLVF) with different $\gamma$ has the same zero sub-level set.
Lemma 8
Proof:
$\ell(\xi(t_{1};t,x,u_{1}(\cdot),\lambda[u_{1}])<0$
for some $t_{1}$. On the other hand, since $V_{\gamma_{2}}(x,t)\geq 0$ and $e^{\gamma_{2}(s-t)}>0$, there exists $\bar{\lambda}$ and for all $u(\cdot)$ s.t.
$\ell(\xi(t_{1};t,x,u(\cdot),\bar{\lambda}[u])\geq 0$
for all $s\in[t,0]$. However, applying $u_{1}(\cdot)$ gives
$\ell(\xi(t_{2};t,x,u_{1}(\cdot),\bar{\lambda}[u_{1}])<0$
for some $t_{2}$, which is a contradiction. Therefore, $V_{\gamma_{1}}(x,t)<0$ immplies $V_{\gamma_{1}}(x,t)\leq 0$. Switching $\gamma_{1}$ and $\gamma_{2}$, we get the same result, and therefore we have $V_{\gamma_{1}}(x,t)\leq 0$ if and only if $V_{\gamma_{1}}(x,t)\leq 0$. With the same process, we could also show that $V^{\infty}_{\gamma_{1}}(x,t)\leq 0$ if and only if $V^{\infty}_{\gamma_{1}}(x,t)\leq 0$. Further, the above inequalities also hold when either $\gamma_{1}$ or $\gamma_{2}$ is 0, and when $\gamma=0$, in the infinite-time horizon, the zero sub-level set is the SRCIS. Therefore, we have proved $\mathcal{Z}_{\gamma}^{\infty}=\mathcal{I}_{m}$.
Note that when $\gamma>0$, the R-CLVF can have a negative value. Consider an initial state in the SRCIS, then $\ell(x)<0$. Then, for its value to be non-negative, $\ell(\xi(s;t,x,u(\cdot),d(\cdot)))$ has to converge to 0 with an exponential rate $\gamma$, which cannot be guaranteed. In fact, in side SRCIS, the R-CLVF is a robust control barrier-value function [29, 36] with respect to the obstacle set defined by $\ell(x)>0$.
The other question to be answered is: given that the R-CLVF with $\gamma$ exists on $\mathcal{D}_{\gamma}$, is $\gamma$ the fastest exponential rate of convergence? Revisting the example used in Remark. 4, we could see that if there exists a control signal for all possible disturbance strategies s.t. the trajectory can reach the $\mathcal{I}_{m}$ in finite time (meaning $\xi(s;t,x,u(\cdot),d(\cdot))\in\mathcal{I}_{m}$ at some $s$), the R-CLVF value at that state will be finite for all $\gamma>0$. Further, if this is the case for all states in an open set, then the R-CLVF exists for all $\gamma>0$ on that open set. This is what happened in that 1D example. On the other hand, if there exists no control signal for all disturbance strategies s.t. the trajectory reaches the $\mathcal{I}_{m}$, we cannot conclude that there exists some $\gamma_{1}>0$ s.t. $\forall\gamma>\gamma_{1}$, the R-CLVF does not exist. In general, we can only conclude the following result:
@@ III-D Impact of the Gamma
From (10) and (14), it can be seen that if $\gamma_{1}>\gamma_{2}$, then $V^{\infty}_{\gamma_{1}}>V^{\infty}_{\gamma_{2}}$. Assume their corresponding domain is $\mathcal{D}_{\gamma_{1}}$ and $\mathcal{D}_{\gamma_{2}}$, we have $\mathcal{D}_{\gamma_{1}}\subset\mathcal{D}_{\gamma_{2}}$. From Theorem 7, we conclude that a larger $\gamma$ corresponds to a faster convergence rate, while a smaller ROES. In other words, the user can trade off between a faster convergence rate and a larger ROES.
With all the results presented, we showed that the R-CLVF is the Lipschitz continuous viscosity solution to the R-CLVF-VI, and satisfies the dynamic programming principle, which provides the theoretical foundation for the numerical computation of the R-CLVF. We also showed that the existence of the R-CLVF is equivalent to the robust exponential stabilizability of the SRCIS of a given system. Further, we showed how the parameter $\gamma$ can affect the R-CLVF. In the next section, we provide a way to synthesize feedback controllers.
III-E R-CLVF-QP
@@
(Feasibility Guaranteed R-CLVF-QP) Given some reference control $u_{r}$, the optimal controller can be synthesized by the following CLVF-QP with guaranteed feasibility $\forall x\in\mathcal{D}_{\gamma}$.
@@
$\hskip 20.00003pt+\max_{d\in\mathcal{D}}D_{x}V_{\gamma}^{\infty}(x)\cdot g_{d}(x)d\leq-\gamma V_{\gamma}^{\infty}(x).$
@@
Note that the QP controller is only point-wise optimal, with respect to “staying close to the reference controller.” It is not optimal w.r.t. the value function. Further, since the R-CLVF is only Lipschitz continuous, its gradient may not be continuous; hence, the QP solution $u=k(x)$ is also not continuous. This may cause the solution of the closed-loop system to lose its uniqueness guarantee [37]. However, such a problem can be solved by considering the sample-and-hold solution as introduced in [38]. The sample-and-hold solution can be viewed as treating the input of the feedback law as a piecewise-continuous (in $s$) input signal, and therefore, the existence and uniqueness can be guaranteed. Further, this type of solution matches with the numerical implementation.
@@ IV-A R-CLVF with Warmstarting
We show this results for three cases: $k(x)=\ell(x)$, $k(x)>\ell(x)$, and $k(x)<\ell(x)$.
@@
The first part is a direct result from Theorem 10. The second part can be proved by contradiction. Assume $x\ \in\bar{\mathcal{D}}_{\gamma}$ but $x\notin{\mathcal{D}_{\gamma}}$. This means $\bar{V}_{\gamma}^{\infty}(x)$ is finite, but ${V_{\gamma}^{\infty}}(x)$ is infinite, which contradicts the first part of this proposition. ∎
The above results are inexact warmstartings, which cannot be used directly in most cases, as we want exact R-CLVF. However, they are vital for the proof of exact warmstarting, which is presented here. We show that for certain initializations, we could recover the exact R-CLVF.
@@
Similar to Propsition 6, If $V_{\gamma}^{\infty}(x)$ exists on $\mathcal{D}_{\gamma}$, then $\bar{V}_{\gamma}^{\infty}(x)\leq{V_{\gamma}^{\infty}}(x)$, and $\mathcal{D}_{\gamma}\subseteq\bar{\mathcal{D}}_{\gamma}$. Combined, we get $\mathcal{D}_{\gamma}=\bar{\mathcal{D}}_{\gamma}$, and $\forall x\in\mathcal{D}_{\gamma}$, $\bar{V}_{\gamma}^{\infty}(x)=V_{\gamma}^{\infty}(x)$.
@@
6:$V^{\infty}(x)\leftarrow$ update_value($f$, $\mathcal{U}$, $\mathcal{D}$, $\Delta$, $\delta t$, $V(x,0)$, $\ell(x)$)
7:$V^{\infty}_{m}\leftarrow\min_{x}V^{\infty}(x)$, $\mathcal{I}_{m}\leftarrow\{V^{\infty}(x)=V^{\infty}_{m}\}$
@@
9:$\ell(x)\leftarrow\ell(x)-V^{\infty}_{m}$, $V(x,t_{0})\leftarrow V^{\infty}(x)-V^{\infty}_{m}$
10:$V_{\gamma}^{\infty}(x)\leftarrow$ update_value($f$, $\mathcal{U}$, $\mathcal{D}$, $\Delta$, $\delta t$, $V(x,0)$, $\ell(x)$)
@@ IV-B R-CLVF with Decomposition
To discuss the R-CLVF with decomposition, we first introduce the self-contained subsystems decomposition.
@@
This is an extension of Lemma 1 of [39], and the proof can be obtained analogously. ∎
@@ V NUMERICAL EXAMPLES
Consider again the system (8), and specify $h(x)=||x||_{\infty}$. We compute the R-CLVF with $\gamma_{1}=0.1$, $\gamma_{2}=0.3$. The results are shown in Fig. 2. It should be noted that for this system, the SRCIS for $\gamma=0.1$ and $\gamma=0.2$ are both $\mathcal{I}_{m}=\{|x|\leq 0.5,|y|\leq 0.5\}$, and ROES $\mathcal{D}_{\text{ROES}}=\{|x|>0.5,|y|<1\}\setminus\mathcal{I}_{m}$.
Figure 2: Top: R-CLVF with $\gamma=0.1$ (left) and $\gamma=0.2$ (right). Bottom left: ROES, SRCIS, and the two optimal trajectories using R-CLVF-QP controller. The ROES and SRCIS for different $\gamma$ are all the same, while the optimal trajectories are different. To see this, first consider a point on the boundary of ROES, $[0.1,1]$, $d$ will make $x$ increase 0.1 to 1, while $u$ cannot decrease $y$. Since the distance is measured by $||x||_{\infty}$, we have $\ell(\xi(s;t,x,u^{*}(\cdot),\lambda[u^{*}]))=1$, $\forall t<0$. Using equation (10) and (14), the value will be infinite. However, for any $|y|<1$, the control can decrease $y$ to 00, and for all $x$, it either goes to 0.5 or -0.5. Note both happen in a finite time horizon. Therefore, using equation (10) and (14), the value will be finite for all $\gamma\geq 0$. Bottom mid: value decay along the two optimal trajectories. All controllers were generated using R-CLVF-QP. With a 151-by-151 grid, the computation time for $\gamma=0.1$ is 215.6s with warmstarting, and 289.7s w/o warmstarting, and 211.5s with warmstarting, and 258.4s w/o warmstarting for $\gamma=0.2$.
@@
Figure 3: Different SRCISs with different $h(x)$. Top left: SRCIS and optimal trajectory with $h(x)=||x||_{2}$. Top right: SRCIS and optimal trajectory with $h(x)=||x||_{Q}$, where $Q=diag[1,1,0]$. Bottom left: the value along the optimal trajectories. All controllers were generated using R-CLVF-QP. With a 51-by-51-by-53 grid, the computation time for $h(x)=||x||_{2}$ is 264s with warmstarting, and 386.6s w/o warmstarting, and 143.4s with warmstarting, and 207.7s w/o warmstarting for $h(x)=||x||_{Q}$.
@@
This 10D system can be decomposed into three subsystems: X-sys with states $[x,v_{x},\theta_{x},\omega_{x}]$, Y-sys with states $[y,v_{y},\theta_{y},\omega_{y}]$, and Z-sys with states $[z,v_{z}]$. It can be verified that all three subsystems have an equilibrium point at the origin. Further, there’s no shared control, disturbance, or state among subsystems. We use $h(x)=||x||_{\infty}$, which satisfies the condition $\ell(x)=\max_{i}\ell_{i}(z_{i})$. The R-CLVF is reconstructed using equation (36).

### v3

@@ III-C R-CLVF implies Exponential Stabilizability
The system can be exponentially stabilized to its smallest robustly control invariant set $\mathcal{I}_{m}$ from $\mathcal{D}_{\gamma}\setminus\text{interior}(\mathcal{I}_{m})$ (or $\mathbb{R}^{n}\setminus\text{interior}(\mathcal{I}_{m}$)), if the R-CLVF exists in $\mathcal{D}_{\gamma}$ (or $\mathbb{R}^{n}$). Further, $\mathcal{D}_{\gamma}\subset\mathcal{D}_{\text{ROES}}$.
@@
Assume the limit in (23) exists in $\mathcal{D}_{\gamma}$. For any initial state $x\in\mathcal{D}_{\gamma}\setminus\mathcal{I}_{m}$, from Proposition 4, because $\mathcal{U}$ is compact, there exists $u^{*}$ s.t. for all $d$,
$D_{x}V_{\gamma}^{\infty}(x)\cdot f(x,u^{*},d)=\dot{V}_{\gamma}^{\infty}\leq-\gamma V^{\infty}_{\gamma}.$
@@
$V_{\gamma}^{\infty}\big(\xi(s;t,x,u^{*}(\cdot),\lambda[u^{*}])\big)\leq e^{-\gamma(s-t)}V_{\gamma}^{\infty}(x),$ (43)
where $u^{*}(\cdot)$ is constructed by $u^{*}$ and $\lambda$ is arbitrary. Consider an arbitrary compact subset of $\mathcal{D}_{\gamma}$, denoted as $\mathcal{S}_{\gamma}$. Define
$\phi(x):=\frac{V_{\gamma}^{\infty}(x)}{\text{dst}(x;\mathcal{I}_{m})}.$
Since $\mathcal{S}_{\gamma}$ is compact and both $V_{\gamma}^{\infty}$ and $dst(x;\mathcal{I}_{m})$ are continuous in $x$, $\phi$ attains its maximum and minimum in $\mathcal{S}_{\gamma}$. This means there exists $c_{1}>0$ and $c_{2}>0$ s.t.
$c_{1}\text{dst}(x;\mathcal{I}_{m})\leq V_{\gamma}^{\infty}(x)\leq c_{2}\text{dst}(x;\mathcal{I}_{m}).$
Combined with (43), we have
$\text{dst}(\xi(s;t,x,u^{*}(\cdot),\lambda[u^{*}]);\mathcal{I}_{m}),$
$\leq$ $\frac{1}{c_{1}}V_{\gamma}^{\infty}(\xi(s;t,x,u^{*}(\cdot),\lambda[u^{*}])),$
$\leq$ $\frac{1}{c_{1}}e^{-\gamma(s-t)}V_{\gamma}^{\infty}(x),$
$\leq$ $\frac{c_{2}}{c_{1}}e^{-\gamma(s-t)}{\text{dst}(x;\mathcal{I}_{m})}.$
Therefore, $\mathcal{S}_{\gamma}$ is a subset of $\mathcal{D}_{\text{ROES}}$. Further, since $\mathcal{S}_{\gamma}$ is an arbitrary compact subset of $\mathcal{D}_{\gamma}$, we conclude that $\mathcal{D}_{\gamma}\subset\mathcal{D}_{\text{ROES}}$.
@@
From (9) and (23), it can be seen that if $\gamma_{1}>\gamma_{2}$, then $V^{\infty}_{\gamma_{1}}\geq V^{\infty}_{\gamma_{2}}$. Assume their corresponding domain is $\mathcal{D}_{\gamma_{1}}$ and $\mathcal{D}_{\gamma_{2}}$, we have $\mathcal{D}_{\gamma_{1}}\subseteq\mathcal{D}_{\gamma_{2}}$. From Theorem 8, we conclude that a larger $\gamma$ corresponds to a faster convergence rate, while a smaller ROES. The user can trade off between a faster convergence rate and a larger ROES.
There is still one question to be answered: given that the R-CLVF with $\gamma$ exists on $\mathcal{D}_{\gamma}$, is $\gamma$ the fastest exponential rate of convergence? Revisting the example used in Remark. 5, we could see that if there exists a control signal for all possible disturbance strategies s.t. the trajectory can reach the $\mathcal{I}_{m}$ in finite time (meaning $\xi(s;t,x,u(\cdot),d(\cdot))\in\mathcal{I}_{m}$ at some $s$), the R-CLVF value at that state will be finite for all $\gamma>0$. Further, if this is the case for all states in an open set, then the R-CLVF exists for all $\gamma>0$ on that open set. This is what happened in that 1D example.
Further, in practice, we typically do not know a priori whether the system is exponentially stabilizable. Consequently, a line search over $\gamma$ may be needed. In this case, it is not necessary to run the computation over the entire discretized grid. Instead, one can construct a smaller grid containing a subset of the SRCIS and perform the computation on this reduced domain, which can significantly decrease the required computational resources.
III-D R-CLVF-QP
@@
(Feasibility Guaranteed R-CLVF-QP) Given $V_{\gamma}^{\infty}$ and a reference control $u_{r}$, the optimal controller can be synthesized by the following CLVF-QP with guaranteed feasibility $\forall x\in\mathcal{D}_{\gamma}$ and $0<\bar{\gamma}\leq\gamma$.
@@
$\hskip 20.00003pt+\max_{d\in\mathcal{D}}D_{x}V_{\gamma}^{\infty}(x)\cdot g_{d}(x)d\leq-\bar{\gamma}V_{\gamma}^{\infty}(x).$
@@
Note that the QP controller is only point-wise optimal, with respect to “staying close to the reference controller.” It is not optimal with respect to the value function. Further, since the R-CLVF is only Lipschitz continuous, its gradient may not be continuous; hence, the QP solution $u=k(x)$ is also not continuous. This may cause the solution of the closed-loop system to lose its uniqueness guarantee [17]. However, such a problem can be solved by considering the sample-and-hold solution as introduced in [16]. The sample-and-hold solution can be viewed as treating the input of the feedback law as a piecewise-continuous (in $s$) input signal, and therefore, the existence and uniqueness can be guaranteed. Further, this type of solution matches the numerical implementation, because we usually do not have an analytic feedback controller.
Note that the R-CLVF is computed for a given rate parameter $\gamma$, whereas in the R-CLVF-QP constraint we enforce a (possibly smaller) rate $\bar{\gamma}$, which relaxes the inequality. With this relaxation, the QP only needs to satisfy a weaker decrease condition, so the optimizer may avoid switching the sign induced by the nonsmooth R-CLVF and might yield a continuous feedback in practice. This makes the QP controller an attractive alternative to the HJ optimal policy, which is typically realized as a bang–bang control.
@@ IV-A R-CLVF with Warmstarting
We show this results for three cases: $\ell(x), k(x)>(x)k$$\ell(x)$, and $k(x)<\ell(x)$.
@@
[…]
The above results concern inexact warm-starting, which cannot be used directly in most cases, since our goal is to recover the exact R-CLVF. Nevertheless, these results are essential for establishing the exact warm-starting presented below. In particular, we show that with appropriate initializations, the warm-started procedure can recover the exact R-CLVF.
@@
Similar to Propsition 5, If $V_{\gamma}^{\infty}(x)$ exists on $\mathcal{D}_{\gamma}$, then $\bar{V}_{\gamma}^{\infty}(x)\leq{V_{\gamma}^{\infty}}(x)$, and $\mathcal{D}_{\gamma}\subseteq\bar{\mathcal{D}}_{\gamma}$. Combined, we get $\mathcal{D}_{\gamma}=\bar{\mathcal{D}}_{\gamma}$, and $\forall x\in\mathcal{D}_{\gamma}$, $\bar{V}_{\gamma}^{\infty}(x)=V_{\gamma}^{\infty}(x)$.
@@
[…]
@@
[…]
@@ IV-B R-CLVF with Decomposition
We first introduce the self-contained subsystems decomposition.
@@
The proof can be obtained analogously following Lemma 1 of [28], and we provide a sketch here. We first prove that $\ell(x)=\max_{i}\ell(z_{i})$ implies
$W_{\gamma}(x,t)=\max_{i}V_{\gamma,i}(z_{i},t)$
by contradiction. More specifically, we assume $W_{\gamma}<V_{\gamma}$, and show that the optimal controls of $V_{\gamma,i}$ can be used to construct a control for the original system. We then use the definition of $W_{\gamma}$ to show that the reconstructed control results a value lower than the optimal control for $V_{\gamma}$, which is a contradiction. The opposite direction can be proved with same procedure. ∎
@@ V NUMERICAL EXAMPLES
Since the numerical value function is obtained on a discrete grid, we use interpolation methods to obtain the value for the states that are not grid points. The R-CLVF-VI may not hold for those states due to the interpolation, and the R-CLVF-QP may lose the feasibility guarantee. To solve this problem, before solving the R-CLVF-QP, we check its feasibility and find the minimal slack variable to make it feasible by solving another optimization problem.
@@
[…]
Figure 2: Top: R-CLVF with $\gamma=0.1$ (left) and $\gamma=0.2$ (right). Bottom left: ROES, SRCIS, and the two optimal trajectories using R-CLVF-QP controller. The ROES and SRCIS for different $\gamma$ are all the same, while the optimal trajectories are different. To see this, first consider a point on the boundary of ROES, $[0.1,1]$, $d$ will make $x$ increase 0.1 to 1, while $u$ cannot decrease $y$. Since the distance is measured by $x||_{\infty}$, we have $\ell(\xi(s;t,x,u^{*}(\cdot),\lambda[u^{*}]))=1$, $\forall t<0$. Using equation (9) and (23), the value will be infinite. However, for any $|y|<1$, the control can decrease $y$ to 0, and for all $x$, it either goes to 0.5 or -0.5. Note both happen in a finite time horizon. Therefore, using equation (9) and (23), the value will be finite for all $\gamma\geq 0$. Bottom mid: value decay along the two optimal trajectories. All controllers were generated using R-CLVF-QP. With a 151-by-151 grid, the computation time for $\gamma=0.1$ is 215.6s with warmstarting, and 289.7s w/o warmstarting, and 211.5s with warmstarting, and 258.4s w/o warmstarting for $\gamma=0.2$.
@@
[…]
@@
[…]